The solubility-product constants, Ksp, at 25 ∘C for two compounds [iron(II) carbonate, FeCO3, and cadmium(II) carbonate, CdCO3] are given by the table Substance Ksp FeCO3 2.10×10−11 CdCO3 1.80×10−14 Part A A solution of Na2CO3 is added dropwise to a solution that contains 1.20×10−2 M Fe2+ and 1.60×10−2 M Cd2+.

What will the concentration of Cd2+ at the moment before Fe2+ begins to precipitate?
Answer #1

FeCO3 -------------> Fe^2+ + CO3^2-

Ksp1   = [Fe^2+][CO3^2-]

[CO3^2-] = Ksp1/[Fe^2+]

CdCO3 --------------> Cd^2+ + CO3^2-

   Ksp2   = [Cd^2+][CO3^2-]

   [Cd^2+] = Ksp2/[CO3^2-]

   [Cd^2+]   = Ksp2/Ksp1/[Fe^2+]

                   = (Ksp2/Ksp1)[Fe^2+]

                    = (1.8*10^-14/2.1*10^-11)(1.2*10^-2)

                    = 0.000857*1.2*10^-2   = 1.0284*10^-5 M